# Differentially Private Federated Learning: A Systematic Review

### v3

@@ 4.2. Differentially Private TFL
Papernot et al. (Papernot et al., 2017) first introduced DP into TFL and proposed the Private Aggregation of Teacher Ensembles (PATE) algorithm. It divides the private dataset into multiple disjoint data batches and trains a “teache” model for each batch. Then, each teacher submits its predictions for the public unlabeled dataset (student) to an aggregator. The aggregator collects and perturbs the predictions of all teachers to ensure that the aggregated labels maintain differential privacy, and sends the perturbed predictions to the students for training. The following year, Papernot et al. (Papernot et al., 2018) optimized the PATE algorithm by using Gaussian mechanism for perturbation and obtaining tighter privacy bounds using RDP, making PATE applicable on a larger scale. Tian et al. (Tian et al., 2022) applied PATE to text generation models. To reduce the noise required by DP for large output spaces in vocabulary size, they dynamically filter out unimportant words to reduce the output space. They also proposed an effective knowledge distillation strategy to reduce the number of queries.
@@ 5. Applications
What we have discussed in the aforementioned parts mostly applied differential privacy FL techniques to classification tasks on image data, as Table 3 shown. In this section, we will describe the more application of differentially private FL from the perspectives of data type and real-world scenarios.
5.1. Application to Data Types.
5.1.1. Graph
Recent advancements in FL applied to graph data focus on developing methodologies that safeguard privacy without sacrificing the utility of graph neural networks (GNNs). The core operation of GNNs is message passing, which makes the learning of a node’s embedding depends not only on the node itself but also on its neighboring nodes. Accurately measuring this dependency level is crucial for determining sensitivity, which determines the noise scale introduced by the DP mechanism (Daigavane et al., 2022). So, the main challenge in applying graph data is controlling the depth and breadth of propagation to limit sensitivity while utilizing the transitivity within graph data, thereby minimizing the impact of DP-introduced noise on model utility.
Currently, methods to address the dependency issue can be divided into two categories. The first category avoids direct measurement of the dependency level by adding noise outside the GNN training phase. Wu et al. proposed LinkTeller, a link-based attack for the vertical federated learning scenario, and implemented link privacy protection by adding Laplacian noise to the graph’s adjacency matrix (Wu et al., 2022). Another approach proposed in (Lin et al., 2022) perturbs decentralized graph data by protecting edge and node features of each user’s adjacency list through randomized response. Under the same setting, Lin et al. suggested perturbing locally trained node embeddings (Wu et al., 2021a). However, these perturbation mechanism results in significant usability loss of the model since the design is not tailored to GNNs. Sajadmanesh et al. proposed extracting the aggregation function from the GNN, performing manual message passing, and adding noise to the aggregated message (Sajadmanesh et al., 2023). This method avoids complex dependency measurement but renders the aggregation function non-learnable, limiting its application to other network architectures. The second category measures the dependency level of nodes and applies gradient perturbation for private GNN training. Daigavane et al. proposed limiting the in-degree of nodes through sampling, measuring the upper bound of the dependency level, and implementing node differential privacy using the DP-SGD paradigm (Dai et al., 2021). According to the theoretical results, the sensitivity remains high, especially as the depth of GNN layers increases.
5.1.2. Bipartite Matrix
@@ 5.1.3. Time Series
5.2. Real-World Implementations
5.2.1. Natural Language Processing
The application of differentially private Federated Learning in natural language processing (NLP) has garnered significant attention, with studies exploring both shallow and large language models. In the realm of shallow networks like recurrent neural networks and long short-term memory (LSTM) models, (McMahan et al., 2017b) and (Kairouz et al., 2021c) propose innovative approaches to training models with sample-level differential privacy while minimizing the impact on predictive accuracy. In the domain of large language models (LLMs), (Wang et al., 2023b) and (Xu et al., 2023b) extend the exploration of differentially private FL, focusing on enhancing privacy-utility trade-offs and implementing practical solutions for real-world applications. Wang et al.  (Wang et al., 2023b) leverages public pre-trained LLMs and introduces a distribution matching algorithm to improve sample efficiency in public training, showcasing a strong privacy-utility trade-off without relying on pre-trained classifiers. Meanwhile, (Xu et al., 2023b) implements DP-FTRL in Google Keyboard (Gboard), ensuring formal differential privacy guarantees without uniform client device sampling. These studies collectively highlight the versatility and effectiveness of differentially private FL techniques in NLP tasks, paving the way for enhanced privacy protection while maintaining utility in language model training across various settings.
@@ 5.2.2. Health Medical
Recent studies have focused on enhancing privacy in the application of FL within the healthcare domain, particularly addressing the concerns around the confidentiality of patient data when leveraging the Internet-of-Medical-Things (IoMT) and distributed healthcare datasets. The research highlighted in (Wu et al., 2021b) introduces the concept of adding artificial noise to IoMT device datasets for user privacy. Similarly, (Malekzadeh et al., 2021) adopts a differentially private stochastic gradient descent approach, combined with secure aggregation through homomorphic encryption, to work on distributed healthcare data, demonstrating its efficacy on a dataset of diabetic retinopathy. Meanwhile, (Choudhury et al., 2019) explores the use of differential privacy in FL to model Electronic Health Records (EHR) across various hospitals, focusing on tasks such as predicting adverse drug reactions and mortality rates, with a significant dataset containing sensitive information like diagnosis and admission records. Further advancements in FL for medical applications aim to address the bandwidth inefficiencies and privacy vulnerabilities associated with the exchange of model updates among a vast network of IoMT devices and federated clients, such as MRI scan machines (Kerkouche et al., 2021; Li et al., 2019; Chen et al., 2021).
@@ 5.2.3. Internet of Things
FL is explored extensively in the IoT recently. Zhao et al. (Zhao et al., 2020b) survey FL applications in mobile edge network comprehensively, including algorithms, applications and potential research problems. Besides, Kong et al. (Kong et al., 2021) propose a collaborative learning framework on the edges for connected vehicles, and it reduces the training time while guaranteeing the prediction accuracy. Lu et al. (Lu et al., 2019) leverage FL to protect the privacy of mobile edge computing and they propose a random distributed update scheme to get rid of the security threats led by a centralized curator, while Pan et al. (Pan et al., 2021b) apply DP to energy harvesting with collaborative and intelligent protection cross the energy side and information side. He et al. (He et al., 2023) propose a LDP scheme to train clustered FL models on heterogeneous IoT data by utilizing dimension reduction methods. Imteaj et al. (Imteaj et al., 2021) studied the problem of how to train distributed machine learning models for resource-constrained IoT devices.
@@ 6.1. Convergence Analysis of Differentially Private HFL
In HFL, A tight upper bound on convergence not only provides theoretical assurance of rapid convergence, but also enables an analysis of the impact of various hyperparameters on the speed of convergence, which can guide parameter tuning or the proposal of new optimization algorithms (Haddadpour and Mahdavi, 2019; Li and Lyu, 2024). However, the convergence analysis of differentially private HFL, compared to HFL, needs to consider the effects of operations such as clipping and adding noise, making it more challenging to obtain a relatively tight upper bound (Zhang et al., 2022). Wei et al. (Wei et al., 2020) introduced the NbAFL algorithm and provided its convergence analysis, while Ling et al. (Ling et al., 2023) offered convergence analysis for DPSGD across multiple iterations, but their assumptions about the objective function were strong. Three studies (Zhang et al., 2022; Cheng et al., 2022; Shi et al., 2023) conducted convergence analysis of their respective CL-DP algorithms under non-convex conditions, addressing the issue of overly strong assumptions on the objective function but still making strong assumptions on gradients, such as Bounded Gradient and Bounded Variance.
@@ 6.2. User-Level DP in HFL
In UL-DP, allowing users to contribute a large number of data samples, even if most contribute only a few, may eventually introduce excessive noise to protect against minority outliers. On the other hand, restricting users to make only small contributions can maintain a lower level of noise but may discard a large amount of surplus data, introducing bias. Some related research has investigated user-level considerations in tasks such as mean estimation, linear regression and experience risk minimisation (Epasto et al., 2020; Liu et al., 2020d; Levy et al., 2021). However, there are currently no articles addressing this issue in the context of federated learning.
@@ 6.3. Privacy Auditing for various DP Models and Neighborhood Levels in FL
6.4. Differentially Private Federated Learning under Large Scale Models
Most of the aforementioned studies have conducted experiments on small datasets and small-scale models. In real industrial scenarios, we face very large datasets and large-scale models. Dealing with extensive text datasets and Large Language Models such as GPT-4 poses significant challenges within the realm of differentially private FL (Wang et al., 2023b; Xu et al., 2023b). The high dimensionality of embeddings means that even small noise can substantially impact the convergence and model performance (Tramer and Boneh, 2020; Yu et al., 2021a). Apart that, the process for methods based DPSGD can be quite time-consuming, as they require clipping for each sample. Therefore, accelerating the training or fine-tuning of DPSGD and enhancing model accuracy in large models become critical issues. Although some efforts have been made in this direction, most current research only focuses on models such as BERT and GPT-2 (Yu et al., 2021b; Shi et al., 2022). Additionally, there remains a significant utility gap between private and non-private models. Moreover, the non-private scenario already necessitates substantial hyperparameter tuning for large-scale models, and the introduction of a privacy perspective further exacerbates this issue, with each tuning attempt adding to the privacy budget (Cattan et al., 2022). This highlights the challenge of efficiently and privately tuning hyperparameters in large models.
@@ 7. Conclusion
In this study, we explored and systematized the differentially private federated learning. We categorized DP models in FL into three major classes, namely DP, LDP, and shuffle model, based on the definitions and guarantees of differential privacy. Further, within DP and shuffle model, we differentiated between SL-DP and CL-DP based on the definition of neighboring datasets. Subsequently, we showed the applications of differentially private federated learning in differential data types and real-world scenarios. Based on these discussions, we provided 5 promising directions for future research. We aim to provide practitioners with an overview of the current technical and application of differential privacy in federated learning, stimulating both foundational and applied research in the future.

### v4

@@ 4.2. Differentially Private TFL
Papernot et al. (Papernot et al., 2017) first introduced DP into TFL and proposed the Private Aggregation of Teacher Ensembles (PATE) algorithm. It divides the private dataset into multiple disjoint data batches and trains a “teache” model for each batch. Then, each teacher submits its predictions for the public unlabeled dataset (student) to an aggregator. The aggregator collects and perturbs the predictions of all teachers to ensure that the aggregated labels maintain differential privacy, and sends the perturbed predictions to the students for training. The following year, Papernot et al. (Papernot et al., 2018) optimized the PATE algorithm by using Gaussian mechanism for perturbation and obtaining tighter privacy bounds using RDP, making PATE applicable on a larger scale. Tian et al. (Tian et al., 2022) applied PATE to text generation models. To reduce the noise required by DP for large output spaces in vocabulary size, they dynamically filter out unimportant words to reduce the output space. They also proposed an effective knowledge distillation strategy to reduce the number of queries. Abourayya et al. (Abourayya et al., 2025) apply consensus methods like majority voting to generate pseudo-labels from hard labels, replacing traditional soft label sharing. Protecting these hard labels via DP enhances privacy protection.
@@ 5. Applications
For the most part, the techniques discussed in previous sections have been demonstrated on relatively standard tasks (e.g. image classification on benchmark datasets, as summarized in Table 2). In this section, we broaden the scope by reviewing how differentially private federated learning is being applied across diverse data types and real-world scenarios. We also provide deeper analysis of the unique challenges and solutions in these application domains.
5.1. Applications across Data Modalities.
5.1.1. Graph Data
Recent advancements in FL applied to graph data focus on developing methodologies that safeguard privacy without sacrificing the utility of graph neural networks (GNNs). The core operation of GNNs is message passing, which makes the learning of a node’s embedding depends not only on the node itself but also on its neighboring nodes. Accurately measuring this dependency level is crucial for determining sensitivity, which determines the noise scale introduced by the DP mechanism (Daigavane et al., 2022). So, the main challenge in applying graph data is controlling the depth and breadth of propagation to limit sensitivity while utilizing the transitivity within graph data, thereby minimizing the impact of DP-introduced noise on model utility. Currently, methods to address the dependency issue can be divided into two categories. The first category avoids direct measurement of the dependency level by adding noise outside the GNN training phase. Wu et al. proposed LinkTeller, a link-based attack for the vertical federated learning scenario, and implemented link privacy protection by adding Laplacian noise to the graph’s adjacency matrix (Wu et al., 2022). Another approach proposed in (Lin et al., 2022) perturbs decentralized graph data by protecting edge and node features of each user’s adjacency list through randomized response. Under the same setting, Lin et al. suggested perturbing locally trained node embeddings (Wu et al., 2021a). However, these perturbation mechanism results in significant usability loss of the model since the design is not tailored to GNNs. Sajadmanesh et al. proposed extracting the aggregation function from the GNN, performing manual message passing, and adding noise to the aggregated message (Sajadmanesh et al., 2023). This method avoids complex dependency measurement but renders the aggregation function non-learnable, limiting its application to other network architectures. The second category measures the dependency level of nodes and applies gradient perturbation for private GNN training. Daigavane et al. proposed limiting the in-degree of nodes through sampling, measuring the upper bound of the dependency level, and implementing node differential privacy using the DP-SGD paradigm (Dai et al., 2021). According to the theoretical results, the sensitivity remains high, especially as the depth of GNN layers increases. proposed a federated graph analysis framework, FEAT, that supports arbitrary downstream graph statistics while protecting individual privacy. Liu et al. (Liu et al., 2024a) proposed a federated graph analysis framework, that supports arbitrary downstream graph statistics while satisfying DP guarantee.
5.1.2. Bipartite Interaction
@@ 5.1.3. Time Series Data
5.2. Applications in Real-World Scenarios
@@ 5.2.1. Health Medical
Recent studies have focused on enhancing privacy in the application of FL within the healthcare domain, particularly addressing the concerns around the confidentiality of patient data when leveraging the Internet-of-Medical-Things (IoMT) and distributed healthcare datasets. The research highlighted in (Wu et al., 2021b) introduces the concept of adding artificial noise to IoMT device datasets for user privacy. Similarly, (Malekzadeh et al., 2021) adopts a differentially private stochastic gradient descent approach, combined with secure aggregation through homomorphic encryption, to work on distributed healthcare data, demonstrating its efficacy on a dataset of diabetic retinopathy. Meanwhile, (Choudhury et al., 2019) explores the use of differential privacy in FL to model Electronic Health Records (EHR) across various hospitals, focusing on tasks such as predicting adverse drug reactions and mortality rates, with a significant dataset containing sensitive information like diagnosis and admission records. Additionally, Zhou et al. (Zhou et al., 2024) extended differentially private federated learning to genomics data analysis tasks (such as cancer classification with bulk RNA-seq), thereby protecting patients’ private information.
@@ 5.2.2. Internet of Things
FL is explored extensively in the IoT recently. Lu et al. (Lu et al., 2019) leverage FL to protect the privacy of mobile edge computing and they propose a random distributed update scheme to get rid of the security threats led by a centralized curator, while Pan et al. (Pan et al., 2021b) apply DP to energy harvesting with collaborative and intelligent protection cross the energy side and information side. He et al. (He et al., 2023) propose a LDP scheme to train clustered FL models on heterogeneous IoT data by utilizing dimension reduction methods. Tao et al. (Tao et al., 2024) combined random sparsification with Gaussian perturbation to achieve DP privacy guarantees and bandwidth-adaptive communication.
5.2.3. Finance and Risk Modeling
Differentially private federated learning has gained traction in finance, powering collaborative credit scoring, fraud detection, and risk analytics without exposing sensitive customer data. For example, the FDML framework introduced by Hu et al. (Hu et al., 2019) enabled banks to jointly train credit risk models on feature-partitioned data. Similarly, the Helen system developed by Zheng et al. (Zheng et al., 2019) demonstrated secure multi-party fraud detection via cryptographic protocols. Nevertheless, privacy threats remain significant. Melis et al. (Melis et al., 2019) showed that federated learning model updates can leak user behavior patterns. Fu et al. (Fu et al., 2022b) further revealed the risk of cross-silo label inference attacks. These findings underscore the core challenges of DP-FL: the injected noise exacerbates the imbalance caused by non-IID financial data. To address these issues, emerging solutions, such as the one proposed by Chowdhury et al. (Roy Chowdhury et al., 2020), combine differential privacy with secure computation to enhance privacy in multi-institutional analytics.
@@ 6.1. Convergence Analysis of Differentially Private HFL
In HFL, A tight upper bound on convergence not only provides theoretical assurance of rapid convergence, but also enables an analysis of the impact of various hyperparameters on the speed of convergence, which can guide parameter tuning or the proposal of new optimization algorithms (Haddadpour and Mahdavi, 2019; Li and Lyu, 2024). However, the convergence analysis of differentially private HFL, compared to HFL, needs to consider the effects of operations such as clipping and adding noise, making it more challenging to obtain a relatively tight upper bound (Zhang et al., 2022). Wei et al. (Wei et al., 2020) introduced the NbAFL algorithm and provided its convergence analysis, while Ling et al. (Ling et al., 2024) offered convergence analysis for DPSGD across multiple iterations, but their assumptions about the objective function were strong. Three studies (Zhang et al., 2022; Cheng et al., 2022; Shi et al., 2023) conducted convergence analysis of their respective CL-DP algorithms under non-convex conditions, addressing the issue of overly strong assumptions on the objective function but still making strong assumptions on gradients, such as Bounded Gradient and Bounded Variance.
@@ 6.2. User-Level DP in HFL
In UL-DP, allowing users to contribute a large number of data samples, even if most contribute only a few, may eventually introduce excessive noise to protect against minority outliers. On the other hand, restricting users to make only small contributions can maintain a lower level of noise but may discard a large amount of surplus data, introducing bias. Some related research has investigated user-level considerations in tasks such as mean estimation, linear regression and experience risk minimization (Epasto et al., 2020; Liu et al., 2020d; Levy et al., 2021; Kato et al., 2024). Although (Kato et al., 2024) represents the pioneering work on UP-DP in FL, this field is still underexplored.
@@ 6.3. Privacy Auditing for various DP Models and Neighborhood Levels in FL
6.4. Differentially Private Federated Learning for Fine-tuning Large Language Models (LLMs)
There is significant potential in using federated learning to fine-tune Large Language Models (LLMs) with differential privacy on local data to preserve privacy (Pan and Wu, 2025). For example, Wang et al.  (Wang et al., 2023c) leverages public pre-trained LLMs and introduces a distribution matching algorithm to improve sample efficiency in public training, showcasing a strong privacy-utility trade-off without relying on pre-trained classifiers. Meanwhile, Xu et al. (Xu et al., 2023b) implements DP-FTRL in Google Keyboard (Gboard), ensuring formal differential privacy guarantees without uniform client device sampling. However, it faces considerable challenges. One major challenge is the substantial communication overhead that arises when applying federated learning to LLM fine-tuning. This is due to the iterative aggregation of model updates from decentralized sources. The frequent transmission of these updates for a complex model like an LLM is extremely costly. While some strategies exist to mitigate these communication challenges, a notable trade-off persists: these methods alleviate communication bottlenecks but cannot provide concrete guarantees for protecting training data privacy.
To address this issue, one can design a communication-efficient federated learning approach with differential privacy guarantees, such as DP-LoRA (Liu et al., 2024c). It leverages the inherent low-rank properties of LLMs by integrating a small set of new weights into the model. By focusing fine-tuning on only these parameters, it drastically reduces communication costs, as only a limited number of updated weights need to be transmitted during distributed training (Liu et al., 2025). Besides these, Differentially private zeroth-order optimization (DP-ZO) has shown promise in fine-tuning LLMs while protecting privacy (Bao et al., 2025), DP-ZO replaces the exact first-order gradient (derived via backpropagation) with a random zero-order approximation based on querying the model’s loss, which significantly accelerates model training. How to apply DP-ZO to federated fine-tuning LLMs is also a significant but challenging problem.
Another important challenge is the trade-off between model utility and privacy budget. DPSGD protects privacy by adding noise to model updates. However, the immense number of parameters in the embedding layers of language models requires a substantial amount of noise. This severely impacts the model’s learning ability, a phenomenon known as the curse of dimensionality (De et al., 2022). A feasible solution is to leverage public data for pre-training and knowledge distillation. One approach is to conduct a large-scale, non-private pre-training on public data, and then fine-tune on the private data. Alternatively, knowledge distillation can be used to distill the knowledge learned from private data into another model, which bypasses the challenge of adding a large amount of noise directly to sensitive, high-dimensional data.
@@ 6.5. Cross-domain differentially private TFL
6.6. Multimodal Federated Learning with Differential Privacy
Federal learning has been widely applied to multimodal settings (e.g., multi-sensor or vision-language tasks) (Che et al., 2023). However, there is a lack of work combining multimodal federated learning with DP to protect complex data combinations. The challenges faced include: complementary cross-modal information can lead to inference attacks, and varying modality dimensions complicate the noise calibration for differential privacy. These issues can severely degrade model utility; for example, DPSGD often fails to converge on high-dimensional multimodal models. Therefore, to mitigate the performance loss caused by differential privacy, effective differentially private federated learning for multimodal data must balance high-dimensional sensitivity and data heterogeneity while preventing cross-modal information leakage.
@@ 7. Conclusion
In this study, we explored and systematized the differentially private federated learning. We categorized DP models in FL into three major classes, namely DP, LDP, and shuffle model, based on the definitions and guarantees of differential privacy. Further, within DP and shuffle model, we differentiated between SL-DP and CL-DP based on the definition of neighboring datasets. Subsequently, we showed the applications of differentially private federated learning in differential data types and real-world scenarios. Based on these discussions, we provided 6 promising directions for future research. We aim to provide practitioners with an overview of the current technical and application of differential privacy in federated learning, stimulating both foundational and applied research in the future.
Appendix A Loss Composition Mechanisms in DP
Definition A.1.
(5) $\begin{split}\begin{aligned} G_{\alpha}(P||Q)=&\operatorname{log}\mathbb{E}_{x\sim P(x)}\bigg[e^{\alpha\operatorname{log}\frac{P(x)}{Q(x)}}\bigg]=\operatorname{log}\mathbb{E}_{x\sim Q(x)}\left[\left(\frac{P(x)}{Q(x)}\right)^{\alpha+1}\right]\end{aligned}\end{split}$
Definition A.2.
Definition A.3.
Based on Rényi divergence, zCDP can be obtained as follows:
Definition A.4.
(8) $D_{\alpha}(\mathcal{A}(D)||\mathcal{A}(D^{\prime}))\leq R\alpha.$
And the following Lemma A.5 defines the standard form for converting $(\alpha,R)$-zCDP to ($\epsilon$, $\delta$)-DP.
Lemma A.5.
[…]
Definition A.6.
(9) $D_{\alpha}(\mathcal{A}(D)||\mathcal{A}(D^{\prime}))\leq R.$
And the following Lemma A.7 defines the standard form for converting $(\alpha,R)$-RDP to ($\epsilon$, $\delta$)-DP.
Lemma A.7.
[…]
Definition A.8.
Definition A.9.
(11) $T(\mathcal{A}(D),\mathcal{A}(D^{\prime}))\geq G_{\mu},$
And the following Lemma A.10 defines the standard form for converting $\mu$-GDP to ($\epsilon$, $\delta$)-DP.
Lemma A.10.
Appendix B Fundamental Mechanisms for Differentially Private FL
Data Types DP LDP
B.1. Perturbation Mechanisms for DP in FL
Gaussian Mechanism (Dwork et al., 2014) is the most popular mechanism for DP in FL. Assume $f:x\to\mathbb{X}$ be a function related to a query. For any $x\in\mathbb{X}$, this mechanism adds noise $n\sim\mathcal{N}(0,\sigma^{2})$ to the $f(x)$ to ensure $(\epsilon,\delta)$-DP. Here $\mathcal{N}(0,\sigma^{2})$ denotes the noise sample from the Gaussian (normal) distribution with mean 0 and variance $\sigma^{2}$, where $\sigma^{2}=\frac{2\Delta f^{2}\log(1.25/\delta)}{\epsilon^{2}}$ and $\Delta f$ is the sensitivity of $f$.
Definition B.1.
Skellam Mechanism (Agarwal et al., 2021) is a response to the issue of the Discrete Gaussian mechanism not being closed under summation. Let $\epsilon>0,\alpha>1$ and $f:x\to\mathbb{X}$ be a function related to a query. For any $x\in\mathbb{X}$, this mechanism adds noise $n\sim\mathrm{Sk}_{0,\mu}$ to $f(x)$ to satisfy $(\alpha,\frac{\alpha\Delta f^{2}}{2\mu}+\min\left(\frac{(2\alpha-1)\Delta f^{2}+6\Delta f}{4\mu^{2}},\frac{3\Delta f}{2\mu}\right))$-RDP. Here $\Delta f$ is the sensitivity of $f$ and $\mathrm{Sk}_{0,\mu}$ denotes sampling from the Skellam distribution with mean 0 and variance $\mu$ as follows.
Definition B.2.
where $I_{v}(x)$ is the modified Bessel function of the first kind.
B.2. Perturbation Mechanisms for LDP in FL
RAPPOR (Erlingsson et al., 2014), also building on RR, incorporates a Bloom filter to facilitate discrete frequency estimation in large domains. Specifically, given a discrete domain $\{1,2,\ldots,d\}$, denoted as $[d]$, and $k$ hash functions $\mathbb{H}=\{H_{1},H_{2},\dots,H_{k}\}$ that map values from $[d]$ to $[k]$. Each user first uses these $k$ hash functions to map their value $v$ into the Bloom filter with $m$ bits and then perturbs each bit in the Bloom filter using RR. To reduce collisions, where two values are hashed to the same set of indices, RAPPOR allocates users into several cohorts, assigning a unique group of hash functions to each cohort. For longitudinal privacy, which protects against attacks during multiple accesses, RAPPOR splits $\epsilon$ into two parts: the first for permanent perturbation via RR and the second for instantaneous perturbation based on the permanent perturbation.